2. A TV set, a stove, a flat iron are connected in series to a 220 V line. The resistance of the TV set is 20 ohms, the stove is 50 ohms and the flat iron is 35 ohms. Find the following;







(a) total resistance



(b) the amount of current flowing in each device



(c) the voltage drop across each device

1
Expert's answer
2022-04-18T08:50:00-0400

(a) total resistance

"R=R_1+R_2+R_3\\
=20+50+35=105\:\Omega"

(b) the amount of current flowing in each device

"I_1=I_2=I_3=\frac{V}{R}\\
=220/105=2.1\:\rm A"

(c) the voltage drop across each device

"V_1=I_1R_1=2.1*20={42\:\rm V}\\
V_2=I_2R_2=2.1*50={105\:\rm V}\\
V_3=I_3R_3=2.1*35={73\:\rm V}"